# What-if Analysis for Business Users: Current Practices and Future Opportunities

### v3

@@ 5.2.1. Tasks
They followed the underneath set of tasks to understand this relationship:
@@
Formulate Hypotheses. Participants often had preconceived hypotheses about how drivers impacted KPI goals, based on their experience, expertise, and external factors. For instance, several participants (P1, P3, P4) shared P8’s expectation that “…expect[ed] newspaper ads to be less effective than internet”.
Translate Hypotheses into Models. Participants refined their hypotheses into simple quantitative models. For example, a marketer (P19) aimed to optimize impression rate (i.e., number of customers reached or tapped by a marketing action) by understanding how marketing channels (e.g., ads, website) contributed to the impression rate KPI by weighing how much each channel was contributing towards one impression. They came up with a rough hypothesis that a potential customer looked at two paid media ads and visited the website, it would result in a complete impression, which was translated into a simple linear model (e.g., 2 $\times$ paid media ads + 1 $\times$ website visit = 1 impression.
Revise Models with Data. Participants frequently adjusted their models by varying parameters based on intuition and trial-and-error. Data played an essential part, and participants relied on existing business intelligence (BI) reports, visualizations, and raw data to refine their models. For example, after reviewing more data and playing with more models, P19 updated their formula to include video completion rate to lead to a more successful impression, changing the model formula to 2 $\times$ ads + 1 $\times$ website visit + 1 $\times$ video completion rate = 1 impression instead. This highlights the need for business users to test various driver-action scenarios (Dimara et al., 2021; Wang et al., 2019) and observe it’s impacts on a given KPI before making strategic decisions.
Explore Multiple Scenarios through WIA Techniques. Overall, all participants explored multiple what-if scenarios. Nearly all participants (20 of 22) manipulated driver or KPI values to observe their effects on other variables in the model, while the remaining participants expressed interest in using such advanced analytics. Examples of the varied WIA techniques discussed in our interviews are shown in Figure 3 and described in order of popularity:
@@
  • Sensitivity Analysis: Maximum participants, 20 out of 22, wanted to observe how changes in driver values (e.g., advertising spend across different channels) affect key performance indicators (KPIs) like sales. For example, P7 wanted to know if “they should do something to the unit price so the return [sales] is better?” Such analysis is commonly referred to as sensitivity analysis, where users explore how altering input variables influences output variables (Rappaport, 1967).
@@
    When using spreadsheets, participants created models or formulas for KPIs, as described in earlier tasks and manually adjusted the driver values to observe how recalculated KPI formulas to avail this technique. They also tweaked formulas to simulate and experiment with different scenarios to explore the impact of changes on various driver combinations, helping guide decisions on next steps. While a few participants (2 out of 22) mentioned existing tools like Google Analytics, Power BI, and Salesforce that offer text-based insights or narratives of few scenarios, they felt these tools were too static. They preferred more dynamic, interactive tools that allow for real-time manipulation of drivers, enabling them to test multiple scenarios and see the immediate impact on KPIs.
@@
  • Driver Importance Analysis: 17 out of 22 participants wanted to determine which drivers (e.g., advertising channels) were most significantly influencing the sales KPI, often known as key influencers or driver importance analysis. Currently, participants relied on manually formulating hypotheses about how drivers related to sales KPIs and performing sensitivity analysis by adjusting driver values from far negative to far positive values to observe changes in sales.
@@
    Participants typically used trial-and-error, adjusting driver values (far negative to far positive) to see which changes had the greatest impact on sales. This manual approach relied heavily on domain knowledge and self-constructed formulas for KPIs, making it difficult to fully understand driver-KPI relationships. Participants acknowledged this limitation, as their approach relied heavily on domain knowledge and guesswork. Although tools like Salesforce Einstein offer a feature that ranks drivers by importance, none of the participants mentioned using it. One participant (P8) noted seeing a similar feature in Power BI’s waterfall chart, which is also seen in Einstein, but found it hard to interpret when quick, actionable insights were needed.
@@
  • Segmentation Analysis: Half the participants expressed interest in segmentation analysis, that involved understanding the needs and behaviors of various drilled-down groups of data such as different “regions” or “user types”, as shared by P10, P14, P18. Participants used this technique to tailor their strategies for distinct data segments. To do this, participants manually segmented data in spreadsheets, copying and pasting slices and dices into separate sheets for individual analysis and comparison.
@@
    No existing tools were mentioned as helpful in this context. While platforms like Tableau and Salesforce Einstein provide text-based insights on certain segments, participants noted them to be limited in scope and often one-off that may not always be actionable or relevant to their specific needs. Additionally, if business requirements for a segment change, these could not be reflected in the underneath data to simulate its impact effectively.
@@
  • Correlation Analysis and Outlier Analysis: 10 of 22 participants expressed the need for learning correlations between different drivers and their KPI to help identify whether, in which direction, and to what extent changes in the drivers are associated with changes in the KPI. To make it easier to digest it, they also wanted to learn a summary of the impacts.
@@
    About three to four participants discussed conducting correlation analysis between drivers to learn the trend of a range of changes made to the drivers on the KPI. They mentioned creating heatmaps to visualize correlations, but also wondered how to justify shifting results to other team members when models were changed.
@@
    Additionally, about three to four participants also mentioned support in detecting anomalies in the data using outlier analysis, but only mentioned currently relying on manual inspections for detecting outliers.
@@
  • Comparison Analysis: Half the number of participants were also eager in comparison analysis where they could actively compare two or more of any of the advanced analyses, but found this to be difficult and confusing currently.
@@
  • Goal Seeking Analysis: Few participants came up with a set of driver actions to trigger the desired changes in their KPI goals. P7, for example shared “…send to different email, …change subject line, …colors, … time” actions to run an email promotional campaign to get “1000 more downloads”. This aligns with the basic concept of goal-seeking analysis.
@@
    One participant (P11) referenced Excel’s GOAL SEEK macro, which predicts the necessary driver values to meet a target KPI but shed light on high expertise levels to effectively use it, while another participant, P21, mentioned not using “it enough or have the time enough to use it… but can already see the value…”.
@@
  • Constrained Analysis: Participants also explicitly shared budget and resource constraints while making decisions. For instance, a participant P2 shared they “obviously have a limited budget about what campaigns to do…” and wanted to know if there were means to include these constraints during analysis, rather than relying on manual data adjustments. This is an example of constrained analysis.
@@
Formulate Initial Actions. After exploring multiple scenarios, participants devised actionable decisions based on their analysis. For example, P19’s revised model led to doubling resources on paid media ads and instructing sales managers to push prospective customers to visit the website and see the complete video explicitly during interactions.
@@ 5.2.2. Tools
All participants shared mainly using spreadsheet applications, primarily Excel for the WIA techniques. Non-technical business users relied on manual methods, explicitly creating formulas and running calculations or using operations like filter and pivot to manually slicing and dicing the raw data in different ways, while technical users employed programming tools and machine learning models (Kandel et al., 2012; Alspaugh et al., 2019; Crisan and Fiore-Gartland, 2021). Only one participant, P10, reported using advanced functions (e.g., LOOKUP, VLOOKUP to retrieve data from large tables) and features (e.g., GOAL SEEK) but highlighted a lack of time to set up and use.
Participants also shared creating charts, but only to observe patterns (e.g., seasonal trends), behaviors (e.g., sales best in the west region), or anomalies (e.g., sudden sales drop every 5 years). Only 5 of 22 visualized their data in dashboards (static and interactive) using a variety of BI tools like Power BI, Salesforce, and Tableau, while 17 others relied on creating Excel charts. Those with access to technical data teams (8 out of 22) or other third-party application tools (e.g., Google Analytics, Marketo, and Pendo) used team or application-generated reports that were generated generally on a daily, weekly, or monthly time-frame. However, those who could not get customized reports easily often struggled with using the BI tools on their own.
@@ 5.2.3. Challenges
We identified several broader challenges faced by business users in leveraging advanced analytics for making data-driven decisions, detailed as follows:
Current Advanced Analytics Use Rudimentary Methods. Challenges with advanced analytics are amplified for business users due to their lack of formal technical training. As discussed, they often rely heavily on spreadsheet tools for performing advanced analytics, which, while “super powerful” (P11), come with drawbacks of “need[ing] a lot of experience with [more complex] macros, etc. which not everyone has[d]” (P11), making even simple analyses inefficient, time-consuming, and prone to errors. For example, P11 recounted frustration dealing with extremely large boilerplate Excel sheets made 20 years ago by some individual and how “absolutely non-replicable the process is and how the person who developed it is still contacted at times to understand some niches”.
Participants also appreciated descriptive and exploratory features in BI tools. For example, visual data exploration capabilities of Tableau were praised, calling them “spectacular” (P5) and “Excel on steroids!” (P4). However, these tools were mainly used for exploring patterns in data lacked capabilities to generate and confirm participant’s hypotheses, necessary to inform data-driven decisions. Further, participants also criticized these tools for being static or minimally interactive, offering only basic data statistics, percentage of goals achieved, and limited drill-down scenarios.
Even for those having advanced analyses features, participants disliked them being cumbersome, noting they “need[ed] to have a developer support to use it… and extensive training” (P4) and describing them as “clunky… [with the] filters, …not very user friendly compared to others” (P5). An employee at one of the BI tool companies also admitted that they “are not proud of not using [their own tool], but it does have a steep learning curve” (P21). Participants also highlighted the lack of state-saving features, limiting their usefulness for iterative analyses. Participants needed more dynamic and interactive tools to simulate various scenarios, test hypotheses, and understand driver-KPI relationships effectively, that allowed for more control and proactive support for advanced analytics techniques.
@@
Business Users Conduct Advanced Analytics without Relying on Technical Users. All participants reported a shortage of analysts in companies, regardless of their size, to handle business users’ requests. They mentioned that data teams also have limited bandwidth and prioritize company-wide initiatives (e.g., “developing models for churn rate analysis” (P4, P5), “customer ordering patterns” (P6), “staff hiring models” (P5, P6)) and automation over their needs of analyses queries. Even when analysts were available, participants expressed frustration due to analysts’ lack of domain knowledge and inefficient communication which made the collaboration “a lot of added work [inconvenience] in itself” (P3). Additionally, under the pressure to make quick decisions, business users often did not have the time or patience to engage with data analysts or scientists, spending mostly 1-4 hours (20 out of 22) and at times 4-6 hours (2 out of 22) daily on analysis themselves without relying on them. These findings highlight a gap between ideal collaborative practices and the real-world challenges of achieving effective communication and understanding and goal alignment between business users and data analysts (Davenport et al., 2010; Provost and Fawcett, 2013).
@@
Unique Challenges Exist Beyond Data Analysis. Before delving into advanced analytics to understand driver-KPI relationships, business users often undertake tasks similar to data analysts (Kandel et al., 2012; Alspaugh et al., 2019; Kandogan et al., 2014), such as acquiring, cleaning, and enriching data and face similar challenges of integrating data from various sources, ensuring data quality, etc. but more aggravated since they are not technical users. When they conduct advanced analytics independently, without the support of technical data scientists or analysts, they face additional challenges. For example, they must balance various business factors, such as budget constraints, economic principles (e.g., the impact of price increases on sales), market conditions (e.g., COVID, interest rates), and social considerations (e.g., employee work-life balance), all while under pressure to make accurate and actionable decisions. Also, participants perform analyses continuously, where they experiment with hypothetical scenarios, refine hypotheses, formulate decision actions, share them with stakeholders, and conduct A/B testing over smaller groups of customers over multiple iterations to execute decision actions that eventually help achieve their goals. Consequently, business users’ analyses to make data-driven decisions are inherently more complex, extending beyond traditional exploratory data analysis.

### v4

@@ 5.2.1. Tasks
They followed the following set of tasks to understand this relationship:
@@
Formulate Hypotheses. Participants often had preconceived hypotheses about how drivers impacted KPI goals, based on their experience, expertise, and external factors. For instance, several participants (P1, P3, P4) shared P8’s expectation that they “…expect[ed] newspaper ads to be less effective than internet”.
Translate Hypotheses into Models. Participants refined their hypotheses into simple quantitative models. For example, a marketer (P19) aimed to optimize impression rate (i.e., number of customers reached or tapped by a marketing action) by understanding how marketing channels (e.g., ads, website) contributed to the impression rate KPI by weighing how much each channel was contributing towards one impression. They came up with a rough hypothesis that if a potential customer looked at two paid media ads and visited the website, it would result in a complete impression, which was translated into a simple linear model (e.g., 2 $\times$ paid media ads + 1 $\times$ website visit = 1 impression.
Revise Models with Data. Participants frequently adjusted their models by varying parameters based on intuition and trial-and-error. Data played an essential part, and participants relied on existing business intelligence (BI) reports, visualizations, and raw data to refine their models. For example, after reviewing more data and playing with more models, P19 updated their formula to include video completion rate to lead to a more successful impression, changing the model formula to 2 $\times$ ads + 1 $\times$ website visit + 1 $\times$ video completion rate = 1 impression instead. This highlights the need for business professionals to test various driver-action scenarios (Dimara et al., 2021; Wang et al., 2019), similar to the learnings of (Dimara et al., 2021; Wang et al., 2019) studies, and observe its impacts on a given KPI before making strategic decisions.
Explore Multiple Scenarios through WIA Techniques. All participants explored multiple what-if scenarios to examine outcomes for various hypotheses and eventually make data-driven decisions. The participants described both WIA techniques that were already employed in their existing workflows, as well as techniques they wanted to use. For example, nearly all participants (20 of 22) manipulated driver or KPI values to observe their effects on other variables in the model in their existing tools, while the remaining participants expressed interest in using such advanced analytics. Figure 2 shows the WIA techniques in the order of popularity among the participants. Below, we describe each of the techniques and the current tools (or lack thereof) used by participants.
@@
[…]
[Shared what-if analysis techniques in study I.]What-if analysis techniques shared by the participants for informing their decisions using data.
@@
  • Sensitivity Analysis: A large number of participants, 20 out of 22, wanted to observe how changes in driver values (e.g., advertising spend across different channels) affect key performance indicators (KPIs) like sales. For example, P7 wanted to know if “they should do something to the unit price so the return [sales] is better?” Such analysis is commonly referred to as sensitivity analysis, where professionals explore how altering input variables influences output variables (Rappaport, 1967).
@@
    All 20 participants performed this using spreadsheets. They created models or formulas for calculating KPIs, as described in earlier tasks and manually adjusted values of various driver columns to observe the recalculated KPI formulas to avail this technique. They also tweaked formulas to simulate and experiment with different scenarios to explore the impact of changes on various driver combinations, helping guide decisions on the next steps. Although they performed this technique, participants complained about manually needing to track the change in their KPIs when alterations to the driver values were made. While a few participants (2 out of 22) mentioned existing tools like Google Analytics, Power BI, and Salesforce that offer text-based insights or narratives of scenarios, they felt these tools were too static. They preferred more dynamic, interactive tools that allow for real-time manipulation of drivers, enabling them to test multiple scenarios and see the immediate impact on KPIs.
@@
  • Driver Importance Analysis: 17 out of 22 participants desired to determine which drivers (e.g., advertising channels) were most significantly influencing the sales KPI, often known as key influencers or driver importance analysis. Currently, they all relied on an indirect way of learning driver importance by manually formulating hypotheses about how drivers related to sales KPIs and performing sensitivity analysis.
@@
    Participants typically used trial-and-error, to edit the driver values from far negative to far positive values and observe corresponding formula calculations of sales that has the greatest impact on sales. This manual table editing approach relied heavily on domain knowledge and self-constructed formulas for KPIs, making it difficult to fully understand driver-KPI relationships. Participants acknowledged this limitation, as their approach relied heavily on domain knowledge and guesswork. Although tools like Salesforce Einstein offer a feature that ranks drivers by importance, none of the participants mentioned using it. One participant (P8) noted seeing a similar feature in Power BI’s waterfall chart, which is also seen in Einstein, but found it hard to interpret when quick, actionable insights were needed.
@@
  • Segmentation Analysis: Half the participants performed segmentation analysis, that involved understanding the needs and behaviors of various drilled-down groups of data such as different “regions” or “user types”, as shared by P10, P14, P18. Participants used this technique to tailor their strategies for distinct data segments. To do this, sorting and filtering techniques were employed to retrieve relevant data, but participants manually segmented data in spreadsheets, duplicating slices and dices of data into separate worksheets for individual analysis and comparison.
@@
    No existing tools were mentioned as being directly helpful in this context. While platforms like Tableau and Salesforce Einstein provide text-based insights on certain segments, participants noted them to be limited in scope, not iterable or controllable (Gathani et al., 2024), and often one-off that may not always be actionable or relevant to their specific needs. Additionally, if business requirements for a segment change, these could not be reflected in the data underneath to simulate its impact effectively.
@@
  • Correlation Analysis and Outlier Analysis: 10 of 22 participants expressed the need for learning correlations between different drivers and their KPI to help identify whether, in which direction, and to what extent changes in the drivers are associated with changes in the KPI. To make it easier to digest, they also wanted to learn a summary of the impacts.
@@
    Of 10 participants, only three to four participants had conducted correlation analysis between drivers to learn the trend of a range of changes made to the drivers and their impact on the KPI. This was performed by creating heatmaps over tables to visualize correlations. However, participants expressed uncertainty about how to justify or communicate the implications of shifting results to other team members when model formulas were changed.
@@
    Additionally, about three to four participants also mentioned detecting anomalies in the data using outlier analysis, where they currently relied on manual inspections (e.g., checking for values that are out of bounds or deviate significantly from expected ranges) for detecting outliers.
@@
  • Comparison Analysis: Half the number of participants also desired the abilities to perform comparison analysis where they could actively compare two or more of any of the advanced analyses, but found this to be difficult and confusing currently. Currently, participants needed to manually duplicate previous tables or analyses to compare them with changes made, resulting in a tedious process and a higher risk of human error.
@@
  • Goal Seeking Analysis: Few participants came up with a set of potential driver actions to trigger the desired changes in their KPI goals. P7, for example shared “…send to different email, …change subject line, …colors, … time” actions to run an email promotional campaign to get “1000 more downloads”. This was currently based off of their experience and domain knowledge, but aligns with the basic concept of goal-seeking analysis.
@@
    Only two participants referenced to an existing tool feature that enables this analysis: Excel’s GOAL SEEK macro, which predicts the necessary driver values to meet a target KPI. However, one participant (P11) commented on the required high level of expertise to effectively use it, while the other participant (P21) mentioned not using “it enough or have the time enough to use it… but can see the value…” if the feature was interactive and easy.
@@
  • Constrained Analysis: Participants also explicitly shared budget and resource constraints that they took into account when making decisions. For instance, a participant P2 shared they “obviously have a limited budget about what campaigns to do…” and wanted to know if there were any features or tools to include these constraints during analysis, rather than relying on manual data adjustments. This is an example of constrained analysis.
@@
Formulate Initial Actions. After exploring multiple scenarios, participants devised actionable decisions based on their analysis. For example, P19’s revised model led to doubling resources on paid media ads and instructing sales managers to push prospective customers to visit the website and see the complete video explicitly.
@@ 5.2.2. Tools
All participants mostly used spreadsheet applications, primarily Excel, for the WIA techniques. In Excel, business professionals relied on manual methods to perform various WIA analyses described above, such as explicitly creating formulas and running calculations to explore multiple hypotheses, editing driver columns and experimenting with formula calculations to analyze sensitivity and driver importance, and using operations like filter and pivot to manually segment raw data for segmentation analysis. They also duplicated data for comparisons. In contrast, as reported from related work, technical professionals employed programming tools and machine learning models for their analyses (Kandel et al., 2012; Alspaugh et al., 2019; Crisan and Fiore-Gartland, 2021). Only one participant, P10, reported using advanced functions and features (e.g., LOOKUP, VLOOKUP to retrieve data, GOAL SEEK) but highlighted a lack of time to set up and use them.
Participants also created charts, but only to observe patterns (e.g., seasonal trends), behaviors (e.g., sales best in the west region), or anomalies (e.g., sudden sales drop every 5 years). Only 5 of 22 visualized their data in dashboards (static or interactive) using a variety of BI tools like Power BI, Salesforce, and Tableau, while 17 others relied on creating Excel charts. Those with access to technical data teams (8 out of 22) or other third-party application tools (e.g., Google Analytics, Marketo, and Pendo) were able to generate charts regularly. However, those who could not get customized reports easily often struggled with using the BI tools on their own.
@@ 5.2.3. Challenges
We identified several broader challenges faced by business professionals in leveraging advanced analytics for making data-driven decisions, detailed as follows:
Advanced Analytics Using Rudimentary Methods. Even though a wide range of existing analytic tools are available, business professionals resorted to using rudimentary methods such as manual table editing, trial-and-error approaches to create formulas, and manual copying and duplicating tables. The participants explained the reliance on these rudimentary methods due to several reasons. First, challenges with advanced analytics are amplified for business professionals due to their lack of formal technical training. While spreadsheets were acknowledged as “super powerful” (P11), the participants often lacked the expertise or “experience with [more complex] macros, etc. which not everyone has[d]” (P11). As a result, they had to use basic or outdated tool features, making even simple analyses inefficient, time-consuming, and prone to errors. For example, P11 recounted frustration dealing with extremely large boilerplate Excel sheets made 20 years ago by some individual and how “absolutely non-replicable the process is and how the person who developed it is still contacted at times to understand some niches”. Such workflows hinder collaboration, reproducibility, and scalability, especially for analyzing larger datasets. Secondly, many state-of-the-art BI tools offer limited prescriptive or predictive analytics capabilities. Participants appreciated the descriptive and exploratory features in current BI tools. For example, the visual data exploration capabilities of Tableau were praised, calling them “spectacular” (P5) and “Excel on steroids!” (P4). However, these tools offer only basic data statistics, percentage of goals achieved, and limited drill-down scenarios, further highlighting the rudimentary nature of these descriptive methods. Finally, even when a tool offers advanced analytic features, its usability and learnability often pose a challenge. Participants noted that many analysis features were “clunky…, …not very user-friendly…” (P5), and they “need[ed] to have developer support to use it… and extensive training” (P4). An employee at one of the BI tool companies also admitted that they “are not proud of not using [their own tool], but it does have a steep learning curve” (P21). Participants also highlighted the lack of state-saving features, limiting their usefulness for iterative analyses. Collectively, these issues illustrate why the participants shied away from advanced features and employed rudimentary methods.
@@
Business Professionals Conduct Advanced Analytics without Relying on Technical Professionals. All participants reported a shortage of analysts in companies, regardless of their size, to handle business professionals’ requests. They mentioned that data teams also have limited bandwidth and prioritize company-wide initiatives (e.g., “developing models for churn rate analysis” (P4, P5), “customer ordering patterns” (P6), “staff hiring models” (P5, P6)) and automation over their needs of analyses queries. Even when analysts were available, participants expressed frustration due to analysts’ lack of domain knowledge and inefficient communication which made the collaboration “a lot of added work [inconvenience] in itself” (P3). Additionally, under the pressure to make quick decisions, business professionals often did not have the time or patience to engage with data analysts or scientists, but spent exorbitant hours, mostly 1-4 hours (20 out of 22) and at times 4-6 hours (2 out of 22) daily on analysis themselves without relying on them. These findings highlight a gap between ideal collaborative practices and the real-world challenges of achieving effective communication and understanding and goal alignment between business professionals and data analysts (Davenport et al., 2010; Provost and Fawcett, 2013).
@@
Unique Challenges Exist Beyond Data Analysis. Before delving into advanced analytics to understand driver-KPI relationships, business professionals often perform tasks similar to those of data analysts (Kandel et al., 2012; Alspaugh et al., 2019), such as acquiring, cleaning, and enriching data. However, they face greater challenges of integrating data from various sources, and ensuring data quality, as they lack technical expertise. When they conduct advanced analytics independently, without the support of technical data scientists or analysts, they face additional challenges. For example, they must balance various business factors, such as budget constraints, economic principles (e.g., the impact of price increase on sales), market conditions (e.g., COVID, interest rates), and social considerations (e.g., employee work-life balance), all while under pressure to make accurate and actionable decisions. Also, participants perform analyses continuously, where they experiment with hypothetical scenarios, refine hypotheses, formulate decision actions, share them with stakeholders, and conduct A/B testing over smaller groups of customers over multiple iterations to execute decision actions that eventually help achieve their goals. Consequently, business professionals’ analyses to make data-driven decisions are inherently more complex, extending beyond traditional exploratory data analysis.
# Task-Based Study Sub-Tasks and Open-Feedback Questionnaire
Sub-Tasks
1 Using driver importance analysis technique, what are the top three and least three drivers that influence the sales KPI?
2 Using sensitivity analysis technique, which advertising channels will you invest in in order of most to least to increase the sales KPI?
3 Using the sensitivity analysis technique, what is the sales achieved if you increase the Demand driver by 5% and the Supply driver by 3%? What can
you say about the demand-supply relationship?
4 Using the sensitivity analysis technique, find the percentage increase in the Unit Price driver needed to achieve the same or more sales if the Demand
and Supply drivers don’t change.
5 Using the summary analysis technique, discuss the trend observed of perturbing the non-advertising channel expenses drivers between a range
of -100% to 700% and step size of 100% on the sales? Do you expect this behavior?
6 Using the summary analysis technique, discuss the trend observed of perturbing the advertising channel expenses.
Which advertising channels would you invest or not invest in to increase the sales of the TV? Do you expect this behavior?
7 Using the goal seeking technique, what is the maximum TV sales you can achieve with the original boundary conditions?
8 What is the sales uplift that you can achieve if you only optimize the advertising channel drivers and the Unit Price driver?
9 How much will the sales increase if you have constraints of \$2-\$115 on the Unit Price driver?
10 Can you achieve a target sales of \$980 million? If yes, what is the driver that needs the maximum uplift? Do you think this is feasible?
Open-Feedback Questionnaire
1 Were these specific tasks relatable to the ones you would try to answer to achieve your goal of increasing sales or any other decisions that you
take in your own work?
2 Do you understand the key driving advertising channels behind sales?
3 Do you understand the relationship between the advertising channel expenses and sales?
4 How open would you be to using such WIA techniques for answering your own business questions?
5 What would be some other advanced analytics techniques that would be helpful for you to achieve your goal?
6 Were you able to reach a decision on investments needed to achieve maximum sales? Explain.
7 How confident are you in the decision you reached? Explain your reasoning.
8 What are your concerns about your inferred decision? Explain your reasoning.
9 What were the challenges you faced to achieve the goal of increasing the sales of the TV company?
10 Do you consider the analysis process you currently used of using the WIA efficient? Explain your reasoning.